# Lifted Forward Planning in Relational Factored Markov Decision Processes with Concurrent Actions

### v1

@@ 3 Relational Factored MDPs
We define rfMDPs based on fMDPs, but include a parfactor model inside the state and action space and transition function. That is, the state variables can now contain PRVs, which are then used in the transition model and the reward function. Also, we have action PRVs, whose value is chosen by the agent. In the following definition we use the term interpretation of a set for a truth-value assignment to each element in the set.
Definition 5 (Relational Factored MDPs).
@@
A relational factored MDP is a tuple $(\mathbf{D},\mathbf{L},\mathbf{X},\mathbf{A},G,\mathbf{R})$. The set $\mathbf{D}$ is a set of constants and the set $\mathbf{L}$ is a set of logvars over $\mathbf{D}$. The set $\mathbf{X}$ is a set of PRVs defined over $\mathbf{L}$. The set of possible interpretations $\mathcal{I}_{\mathbf{X}}$ for the groundings of the set $\mathbf{X}$ defines the state space. The set $\mathbf{A}$ is a set of action PRVs. A prafctor model $G$ over $\mathbf{A}$ and $\mathbf{X}$ represents the the transition function $T:\mathcal{I}_{\mathbf{X}}\times\mathcal{I}_{\mathbf{A}}\times\mathcal{I}_{\mathbf{X}}\mapsto\mathbb{R}_{0}^{+}$, with the set $\mathcal{I}_{\mathbf{A}}$ of possible interpretations of the groundings of $\mathbf{A}$, and specifies the transition probability given an action and a previous state. The set $\mathbf{R}$ contains parameterized local reward functions $R_{i}:\bigtimes_{j}\mathcal{R}(B_{i,j})\mapsto\mathbb{R}$, defined over PRVs $B_{i,j}$. The reward function $R$ is decomposed as a sum over $\mathbf{R}$.
@@ 3.1 Parameterizing Actions
Definition 6 (Action PRV).
An action PRV $A$ is a Boolean-valued PRV. A concrete action is a set of events, in which each grounding of $A$ receives an assignment $a\in\mathcal{R}(A)$.
Action PRVs allow for a more general action setting. In our example, the mayor can restrict multiple persons from travelling at once:
Example 4 (Action PRV).
The action PRV $Restrict(M)$ models the possible travel bans on the population of the town. For a concrete action, the mayor has to specify, possibly using constraints on parfactors, on which persons $Restrict$ should be applied on.
@@
3.2 Parameterized Local Reward Functions
The reward function in fMDPs maps from the (joint) state to the reward of the state. For evaluating the reward function, we thus have to construct the joint state and cannot exploit the factorization, breaking our aim of efficient representation. To further use our efficient representation, we introduce a decomposable reward function: We assume that the reward function is factored as $R=\sum_{i}R_{i}$, with local reward functions $R_{i}$ with scope restricted to a subset of the state variables. As we have indistinguishable state variables, we reduce redundant computations by using representatives in the reward functions analogue to a parfactor, but using a sum instead of a product:
Definition 7.
A local reward function $R_{i}:\bigtimes_{j}\mathcal{R}(B_{i,j})\mapsto\mathbb{R}$ is defined over a sequence of PRVs $(B_{i,j})_{j}$. The semantics of a single parameterized local reward function $R_{i}$ is defined as the sum $\sum_{z}R_{i}(z)$ over the interpretations $z\in\bigtimes_{j}\mathcal{R}(B_{i,j})$ in the current state of all groundings of $R_{i}$.
@@ 3.2 Parameterized Local Reward Functions
The parameterized local reward functions for Example 3 are $R_{1}(Sick(M))$, evaluating to $-1$ ($1$) for each person (not) being sick, and $R_{2}(Travel(M))$, evaluating to $2$ for each person travelling. If five persons are sick, three are not sick and four people are travelling, the total reward is $-5+3+8=6$.
Summarizing, we have defined rfMDPs as fMDPs using a parfactor model for states, actions and transitions and adapted the reward function for taking PRVs into account as well. In the next section, we propose a compact state representation for such rfMDPs to be used in Foreplan exploiting indistinguishable objects.
4 Foreplan: Efficient Forward Planning
@@
In this section, we propose Foreplan, our exact forward relational planner for rfMDPs. The input to Foreplan is an rfMDP. The output is a value function, which induces a policy. Foreplan computes the value function by finding an efficient state representation for the rfMDP and then running a linear program based on the state representation to calculate the value function. We first describe how Foreplan uses the state representation for rfMDPs and how it obtains the state representation. Afterwards, we outline how Foreplan computes a value function based on the found state representation.
4.1 Obtaining an Efficient State Representation
@@
Foreplan needs to encode the current state compactly to efficiently reason about indistinguishable variables. Thus, in this subsection, we describe how Foreplan treats indistinguishable objects in rfMDPs. Namely, Foreplan can tame the information to store: Foreplan does not need to keep track of objects that could be differentiated by their history. Rather, with each action and new time step, the history is swept away and the objects remain indistinguishable because of the first-order Markov assumption. The basic idea of our state representation is inspired by Counting Random Variables (CRVs) [30] and we extend them for use in out state representation. It is sufficient to count the number of occurrences of each possible truth-value assignment to the groundings of the input PRVs of a parfactor in a histogram. The input PRVs are the ones representing the current state. Counting only the input PRVs is sufficient, because all possible next states are iterated separately, which we describe in the next subsection. Focusing only on the counts for the groups enables Foreplan to use a much simpler state space representation, namely the set of possible histograms. We now describe in more detail how to count the assignments.
@@
Counting the assignments of each PRV separately is insufficient, as PRVs can be defined over the same logvars and thus interfere with each other. However, the parfactors can be evaluated separately since, in Equation 2, we have the full current and next state available. Thus, it is sufficient to count PRVs together if they share a logvar and occur in a parfactor together. To obtain the representation and quantify its complexity, we define the relational cost graph:
Definition 8 (Relational Cost Graph).
@@
The relational cost graph of a parfactor model of an rfMDP has a vertex for each PRV in the current state. Two vertices are connected by an edge if and only if the PRVs associated with these two vertices share a logvar and occur together in a parfactor or a parameterized local reward function. We denote the number of (maximal) cliques by $c$ and the size of the largest clique by $w$.
We note that $c$ and $w$ are both bound by the number of PRVs, although this bound is relatively loose. The key insight now is that (maximal) cliques in the relational cost graph correspond to sets of PRVs that Foreplan needs to count together as they interfere with each other. Let us take a look at the relational cost graph of Example 3. For a more complex example, we refer to Appendix A.1.
Example 6.
@@
The relational cost graph for Example 3 consists of three isolated vertices corresponding to $Sick(M)$, $Travel(M)$ and $Epidemic$. The first two do not occur together in a parfactor or local reward function and the first and the last does not share a logvar with one of the first two.
Since the cliques describe which PRVs need to be counted together, the state representation is now a set of such countings, stored in a histogram each. We extend the definition of CRVs to more than one logvar based on the definition of CRVs by Taghipour [30]:
Definition 9 (Extended Counting Random Variable).
@@
If no restrictions apply, we omit $C$, and $n=\prod_{L\in\mathbf{L}}\left\lvert\mathcal{D}(L)\right\rvert$, where $\mathbf{L}$ is the set of logvars of the PRVs $B_{i}$. The CRV corresponding to a clique gives us the number of occurrences for each possible instantiation of the PRVs in that clique. We give a small example for a CRV using an additional PRV $Car(M)$ for illustrative purposes:
Example 7.
Suppose we have two PRVs, $Travel(M)$ and $Car(M)$. A possible state for the CRV #$\#[Travel(M),Car(M)]$ is $\{(tt,2),(tf,1),(ft,3),(ff,2)\}$. In this state, there are two persons for which $Travel$ and $Car$ are both true.
Example 7 illustrates that we need only four buckets regardless of the domain size of $M$. We now formalize the state representation:
Definition 10 (State Representation).
@@
Example 8.
As the vertices are not connected in the relational cost graph, the state representation is (#$(\#[Sick(M)],\#[Travel(M)],Epidemic)$.
We prove that our state representation exactly covers $\mathbf{S}$:
Theorem 11.
The representation in Definition 10 is correct.
@@
Given groundings for the state PRVs, we derive the histograms for the CRVs by counting the assignments for each parfactor. Given a representation as defined in Definition 10, we reconstruct, for each parfactor, the groundings of the PRVs by extracting the counts from the CRVs and instantiating the respective parfactors. We provide a full proof in Appendix A.2. ∎
@@
Example 9.
The action $Restrict$ works on the PRV $Travel(M)$. Thus, the mayor needs to specify how many persons of those (not) travelling are (not) allowed to travel. The action $Restrict$ is therefore defined over #$\#[Travel(M),Restrict(M)]$. A concrete action is, e.g., $a=\{(tt,3),(ft,2)\}$. The action $a$ does not need to specify the counts of people no travel ban is imposed on ($tf$, $ff$), as these are determined by $a$ and the current state.
The mayor no longer needs to specify individual persons, but rather the number of persons (not) travelling, which are restricted from travelling. It is irrelevant on which exact persons the action is performed. With this action representation, we reduce the action space from exponential to polynomial, which we prove in Theorem 13 in the next section.
@@
4.2 Computing a Value Function
@@
We briefly describe how to evaluate the transition probability $P(s^{\prime}\mid s,a)$. Since we have full evidence provided, Foreplan evaluates each state CRV $S_{i}^{\prime}$ separately. For a fixed state CRV $S_{i}^{\prime}$, the value $s_{i}^{\prime}$ is fixed since the whole state space is iterated. For evaluating $P(s_{i}^{\prime}\mid s,a)$, Foreplan iterates over all possible assignment transitions, e.g., the number of people getting sick and healthy, and calculating the transition probability for a single assignment transition by using the parfactor as a lookup-table for the probabilities for the transition and using the multinomial coefficient for counting how many times the assignment transition is applicable. We provide a more detailed description and an example in Appendix A.3.
@@
5 Complexity Analysis of Foreplan
Having outlined Foreplan, we analyze the complexity of Foreplan in this section. We start by quantify the state representation and using the complexity of the state representation to derive the runtime complexity of Foreplan.
We derive the following theorem about the size of the state representation from Definition 10.
Theorem 12.
The state representation is in $\mathcal{O}(c\times 2^{w})$.

### v2

@@ 3. Relational Factored MDPs
We define rfMDPs based on fMDPs, but include a parfactor model inside the state and action space and transition function. That is, the state variables can now contain PRVs, which are then used in the transition model and the reward function. Also, we have action PRVs, whose value is chosen by the agent. Basically, we keep the same semantics as in Definition 2.1 and only change the representation to exploit indistinguishable variables. In the following definition we use the term interpretation of a set for a truth-value assignment to each element in the set.
[…]
@@
A relational factored MDP is a tuple $(\mathbf{D},\mathbf{L},\mathbf{X},\mathbf{A},G,\mathbf{R})$. The set $\mathbf{D}$ is a set of constants and the set $\mathbf{L}$ is a set of logvars over $\mathbf{D}$. The set $\mathbf{X}$ is a set of PRVs defined over $\mathbf{L}$. The set of possible interpretations $\mathcal{I}_{\mathbf{X}}$ for the groundings of the set $\mathbf{X}$ defines the state space, i.e., all possible assignments. The set $\mathbf{A}$ is a set of action PRVs. A parfactor model $G$ over $\mathbf{A}$ and $\mathbf{X}$ represents the transition function $T:\mathcal{I}_{\mathbf{X}}\times\mathcal{I}_{\mathbf{A}}\times\mathcal{I}_{\mathbf{X}}\to\mathbb{R}_{0}^{+}$, with the set $\mathcal{I}_{\mathbf{A}}$ of possible interpretations of the groundings of $\mathbf{A}$, and specifies the transition probability given an action and a previous state. The set $\mathbf{R}$ contains parameterized local reward functions $R_{i}:\bigtimes_{j}\mathcal{R}(B_{i,j})\to\mathbb{R}$, defined over all PRVs $B_{i,j}$, which contribute to the local reward $R_{i}$. The reward function $R$ is decomposed as a sum over $\mathbf{R}$.
Let us generalize Example 2.1 to an arbitrary number of persons behaving in the same way as the agent in Example 2.1:
[…]
There is a set $\mathbf{D}$ of persons living in a small town, represented by the logvar $M$ with $\mathcal{D}(M)=\mathbf{D}$. Each person can be sick or healthy, leading to the PRV $Sick(M)$. The government gets a reward of $1$ for each healthy person and a reward of 1-1 for each sick person. To combat an epidemic, the government can impose travel bans on persons, resulting in the action $Restrict(M)$ to impose a travel ban on a subset of persons. Moreover, each person can travel or not, leading to the PRV $Travel(M)$. The government gets a reward of $2$ for each person travelling. The PRV $Epidemic$ is influenced by the number of people travelling and influences the healthiness of each person. Figure 1 shows the grounded transition model and Figure 2 shows the lifted transition model for this example.
[…]
Figure 1. Grounded transition model for Example 3. We assume that the population of the town consists of two citizens $a$ and $b$. For abbreviation, we use only the first letter(s) for each symbol. The factors $f_{1}$ and $\tilde{f}_{1}$ as well as $f_{2}$ and $\tilde{f}_{2}$ could be combined into a parfactor. Figure 2 shows the lifted version.
[…]
Grounded graphical model for the epidemic example: The factor f_1 links Travel(a) and Restrict(a) to Travel’(a), analogously tilde f_1 for b, f_2 links Sick(a) and Epidemic to Sick’(a), analogously tilde f_2 for b, and f_3 links Travel(a) and Travel(b) to Epidemic’.
[…]
Figure 2. Lifted representation of the transition model for Example 3. We abbreviate by using only the first letter(s) for each symbol.
[…]
Since an action can be applied to each person concurrently, the amount of possible actions is exponential due to the power set, i.e., all possible combinations for all persons. The action PRVs help Foreplan to reduce the exponential blow-up by exploiting the indistinguishability.
With Definition 3, we require the rewards to be represented as a sum of local reward functions, and not as some arbitrary function. Actually, this is too restrictive: It is sufficient that an operation running over same behaving variables has a hyperoperation repeating the operation multiple times. For a sum, the hyperoperation is multiplication. Also, multiple terms of such operations can be linked together in any way. In Appendix A.1, we give examples for hyperoperations. For simplicity, we use the definition as it is throughout the paper.
@@ 3.1. Parameterizing Actions
[…]
An action PRV $A$ is a Boolean-valued PRV. A concrete action is a set of events, i.e., each grounding of $A$ receives an assignment $a\in\mathcal{R}(A)$.
Action PRVs allow for a more general action setting. When writing action, we refer to Definition 3.1, i.e., concurrent actions. In our example, the mayor can restrict multiple persons from travelling at once:
[…]
The action PRV $Restrict(M)$ models the possible travel bans on the population of the town. For a concrete action, the mayor has to specify which persons are restricted from travelling. When restricting the persons $a$, $b$ and $f$, the mayor specifies $Restrict(M)=true$ with constraint $(M,\{a,b,e\})$ and $Restrict(M)=false$ with constraint $(M,\{c,d,f,g,h\})$.
When the persons $a$ to $h$ are indistinguishable, it is irrelevant which permutation is used for the constraints, only the amounts are relevant, which we exploit in Foreplan.
@@
3.2. Parameterized Local Reward Functions
The reward function in fMDPs maps from the (joint) state to the reward of the state. For evaluating the reward function, we thus have to construct the joint state and cannot exploit the factorization, breaking our aim of compact representation. To further use our compact representation, we introduce a decomposable reward function: For simplicity, we assume that the reward function is factored as $R=\sum_{i}R_{i}$, with local reward functions $R_{i}$ with scope restricted to a subset of the state variables. Other operations are possible too, as long as they have a hyperoperation, as already mentioned. As we have indistinguishable state variables, we reduce redundant computations by using representatives in the reward functions analogue to a parfactor, but using a sum instead of a product:
[…]
A local reward function $R_{i}:\bigtimes_{j}\mathcal{R}(B_{i,j})\to\mathbb{R}$ is defined over all PRVs $B_{i,j}$, which contribute to the local reward $R_{i}$. The semantics of a single parameterized local reward function $R_{i}$ is defined as the sum $\sum_{z}R_{i}(z)$ over the interpretations $z\in\bigtimes_{j}\mathcal{R}(B_{i,j})$ in the current state of all groundings of $R_{i}$.
@@ 3.2. Parameterized Local Reward Functions
The parameterized local reward functions for Example 3 are $R_{1}(Sick(M))$, evaluating to 1-1 (11) for each person (not) being sick, and $R_{2}(Travel(M))$, evaluating to $2$ for each person travelling. If five persons are sick, three are not sick and four people are travelling, the total reward is 5+3+8=6-5+3+8=6.
With rfMDPs, we can efficiently represent the action and state space as well as the transition function for numerous indistinguishable objects. Further, we have a compact representation for concurrent actions. In the next section, we propose Foreplan, our main contribution, which uses our representation to solve planning in rfMDPs in time polynomial in the number of objects.
4. Foreplan: Efficient Forward Planning
@@
In this section, we propose Foreplan, our exact forward relational planner for rfMDPs. The input to Foreplan is an rfMDP. The output is a value function, which induces a policy. Foreplan computes the value function by using a compact state representation for the rfMDP to exploit the indistinguishability of the objects and then running a linear program based on the state representation to calculate the value function.
The first-order representation in rfMDP contains dependencies between PRVs. For using the representation in computation, we first need to recognize the dependencies and find a suitable representation of them to compute the value function. We first describe how Foreplan recognizes the dependencies and identifies a suitable representation, Afterwards, we outline how Foreplan uses the identified representation to compute a value function.
4.1. Obtaining a Compact State Representation
@@
Foreplan needs to encode the current state compactly to efficiently reason about indistinguishable variables. Thus, in this subsection, we describe how Foreplan treats indistinguishable objects for computations using rfMDPs. The representation in rfMDPs does not consider dependencies between PRVs. But, Foreplan can tame the information to store: Foreplan does not need to keep track of objects that could be differentiated by their history. Rather, with each action and new time step, the history is swept away and the objects remain indistinguishable because of the first-order Markov assumption. Thus, the basic idea is to count the objects for which the PRVs are true or false, which is inspired by Counting Random Variables (CRVs) crv. When evaluating a parfactor, the joint assignment to the PRVs is needed, rather than individual counts for each PRV. It is sufficient to count the number of occurrences of each possible truth-value assignment to the groundings of the input PRVs of a parfactor in a histogram. The input PRVs of a parfactor are the ones representing the current state. Counting only the input PRVs is sufficient, because all possible next states are iterated separately, which we describe in the next subsection. But some parfactors may be defined over PRVs defined over different logvars, like $f_{3}$ in Figure 2 is defined over $Travel(M)$ and $Epidemic$. Given the counts for $Travel(M)$ and $Epidemic$ separately is enough, since these PRVs do not depend on each other. But how does Foreplan recognize which PRVs depend on which PRVs? For answering this question, we propose the relational cost graph to identify the dependencies between PRVs and identify the groups that need to be counted jointly. Focusing only on the counts for the groups enables Foreplan to use a much simpler state space representation, namely the set of possible histograms, compared to the grounded state space representation, which is the Cartesian product over all state variables. We now describe in more detail how to count the assignments.
@@
Counting the assignments of each PRV separately is insufficient, as PRVs can be defined over the same logvars and thus interfere with each other. However, the parfactors can be evaluated separately since, in Equation 2, we have the full current and next state available. Thus, it is sufficient to count PRVs together if they share a logvar and occur in a parfactor together. To obtain the representation and later on quantify its complexity, we define the relational cost graph:
[…]
@@
The relational cost graph of an rfMDP has a vertex for each PRV in the current state. Two vertices are connected by an edge if and only if the PRVs associated with these two vertices share a logvar and occur together in a parfactor as input PRVs or a parameterized local reward function. We denote the number of (maximal) cliques by $c$ and the size of the largest clique by $w$.
We give a more formal definition of the relational cost graph in Appendix A.2. Let us take a look at the relational cost graph of Example 3. For a more complex example, we refer to Appendix A.3.
[…]
@@
The relational cost graph for Example 3 consists of three isolated vertices corresponding to $Sick(M)$, $Travel(M)$ and $Epidemic$. The first two do not occur together in a parfactor or local reward function and the last does not share a logvar with one of the first two. Figure 3 shows the visual representation of the relational cost graph.
Si(M)Si(M)Tr(M)Tr(M)EpiEpi
Figure 3. Visual representation of the relational cost graph in Example 3. We abbreviate using the first letter(s).
[…]
The relational cost graph for the epidemic example consists of three isolated vertices corresponding to Sick(M), Travel(M) and Epidemic.
The key insight now is that (maximal) cliques in the relational cost graph correspond to sets of PRVs that Foreplan needs to count together as they interfere with each other. But when having more than one logvar in a PRV, basic CRVs crv only allow to count one of them. To use our representation in computations, we extend CRVs to allow for more than one logvar, extending the definition by crv:
[…]
@@
If no restrictions apply, we omit $C$, and $n=\prod_{L\in\mathbf{L}}\left\lvert\mathcal{D}(L)\right\rvert$, where $\mathbf{L}$ is the set of logvars of the PRVs $B_{i}$. We do not define the operations to manipulate CRVs as we do not need them in this paper. The CRV corresponding to a clique gives us the number of occurrences for each possible instantiation of the PRVs in that clique. We give a small example for a CRV using an additional PRV $Friends(M,Y)$, with $M$ and $Y$ having the same domain, for illustrative purposes:
[…]
To have a PRV with two logvars, assume we have three PRVs, $Travel(M)$, $Friends(M,Y)$ and $Sick(Y)$ with $\left\lvert\mathcal{D(M)}\right\rvert=\left\lvert\mathcal{D}(Y)\right\rvert=8$. A possible state for the CRV #[Travel(M)\#[Travel(M), $Friends(M$, $Y)$, $Sick(Y)]$ is $\{(ttt,6)$, $(ttf,30)$, $(tft,4)$, $(tff,0)$, $(ftt,6)$, $(ftf,0)$, $(fft,0)$, (fff,18)}(fff,18)\} and we have $n=64$. The first histogram tells, that in this state, there are six tuples for which $Travel(a)=t$, $Friends(a,b)=t$, and $Sick(b)=t$ holds.
Example 4.1 illustrates that we need only eight buckets in a histogram regardless of the domain size of $M$ and $Y$. The number of buckets depends on the range values of the PRVs and not on the domain sizes of them. We now formalize the state representation, which is basically a set of CRVs per (maximal) clique in the relational cost graph:
[…]
@@
[…]
As the vertices are not connected in the relational cost graph, so the state representation is (#[Sick(M)](\#[Sick(M)], #[Travel(M)]\#[Travel(M)], $Epidemic)$.
The relational cost graph in Example 4.1 tells us that we can count the number of healthy persons separately from the number of travelling persons. In particular, we do not store which sick persons are travelling, since we do not need this information. We prove that our state representation is correct, i.e., retains the same semantics, namely exactly representing $\mathbf{S}$: {theorem} The representation in Definition 4.1 is correct.
@@
Given groundings for the state PRVs, we derive the histograms for the CRVs by counting the assignments for each parfactor. Given a representation as defined in Definition 4.1, we reconstruct, for each parfactor, the groundings of the PRVs by extracting the counts from the CRVs and instantiating the respective parfactors. We provide a full proof in Appendix A.4. ∎
@@
[…]
The action $Restrict(M)$ uses the PRV $Travel(M)$. Thus, the mayor needs to specify how many persons of those (not) travelling are (not) allowed to travel. Suppose that in the current state five out of eight persons are travelling and two out of eight persons are sick. The action $Restrict(M)$ is defined over #[Travel(M),Restrict(M)]\#[Travel(M),Restrict(M)]. A concrete action is, e.g., $a=\{(tt,3)$, (ft,2)}(ft,2)\}, which means that three persons currently travelling are restricted from travelling and two persons not travelling. The action $a$ does not need to specify the counts of people no travel ban is imposed on ($tf$, ffff), as these are determined by $a$ and the current state.
The mayor no longer needs to specify individual persons (c.f. Example 3.1), but rather the number of persons (not) travelling, which are restricted from travelling. It is irrelevant on which exact persons the action is performed. With this action representation, we reduce the action space from exponential to polynomial, which we prove in Theorem 5 in the next section.
@@
4.2. Computing a Value Function
@@
For instantiating the linear program in Equation 2, it is left to describe how Foreplan calculates the transition probability. Since we have full evidence provided, Foreplan evaluates each state CRV $S_{i}^{\prime}$ separately. For a fixed state CRV $S_{i}^{\prime}$, the value $s_{i}^{\prime}$ is fixed since the whole state space is iterated. For evaluating $P(s_{i}^{\prime}\mid s,a)$, Foreplan iterates over all possible assignment transitions, e.g., from each bucket in the histogram $s$ to each bucket in the histogram $s^{\prime}$. In our epidemic example, this is, e.g., the number of people getting sick and healthy. For each possible assignment transition, Foreplan calculates the transition probability: The probability for a representative transition is given by the involved parfactors, i.e., the parfactors defined over the PRVs mentioned in $S_{i}^{\prime}$. But this representative probability has to be weighted by how many times this assignment transition is applicable using the multinomial coefficient. Take for example the state in which five persons are sick. When three sick persons get healthy, this assignment transition has weight $5\choose 3$ as any three of the five persons can get healthy. The final probability of $P(s_{i}^{\prime}\mid s,a)=\sum_{t}w_{t}p_{t}$ for a fixed CRV is the sum over all possible assignment transitions $t$, each with representative probability $p_{t}$ and weight $w_{t}$. Since the state space is factored, the full transition probability $P(s^{\prime}\mid s,a)$ is factored as $P(s^{\prime}\mid s,a)=\prod_{i}P(s_{i}^{\prime}\mid s_{i},a)$. We provide a more detailed description and an example in Appendix A.5.
Let us have a closer look at the actual creation of Foreplan’s linear program: For each state and action combination, a constraint is added. Thus, Foreplan can take both, state and action, into account for generating only the necessary constraints. When an initial state is given, the reachable state space can be pruned, and Foreplan limits all computations to the pruned search space, that is why we are calling Foreplan a forward planner. Moreover, additional checks like mutual exclusion, capacities, or other constraints can be added, either on PRV level or on search space level. We make this more concrete with an example: Assume we have four persons travelling. Then, since the action uses the lifted state representation, the mayor can only restrict up to four persons, which currently travel, from travelling. The generation of constraints for the assumed state only iterates over the possible actions. Future work includes an analysis of what types of additional applicability checks are efficient implementable.
@@
5. Complexity Analysis of Foreplan
Having outlined Foreplan, we analyze the complexity of Foreplan in this section. We start by quantifying the state representation and using the complexity of the state representation to derive the runtime complexity of Foreplan.
We derive the following theorem about the size of the state representation from Definition 4.1. {theorem} The state representation is in $\mathcal{O}(c\times 2^{w})$.